OpenStudy (anonymous):

h(t)=-16t^2+32t+3 1. a) What are the coordinates of the vertex of the parabola? b) What do the coordinates of the vertex of the parabola mean in the context of this problem? 2. How high off the ground will the ball be after 2 seconds? 3. Why is the height the same at .5 seconds and 1.5 seconds and what is the height? 4. a) How long is the ball in flight? Round answer to nearest hundredth. b) Which formula, process, etc. did you use to solve this problem?

OpenStudy (anonymous):

to get the vertex of the parabola, let v=-b/(2a) = -32/(2*(-16)) = 1. it is at this time, t=1 where the vertex occurs. to find height at this time, plug into the equation for h.

OpenStudy (anonymous):

thats what my teacher said but i just kept gettin confused

OpenStudy (amistre64):

h(t)=-16t^2+32t+3 h(t=1)= -16(1)^2+32(1)+3 = -16+32+3 = ??

OpenStudy (amistre64):

what are you gettiing confused about?

OpenStudy (anonymous):

i get lost when i have to distribute and do -b/(2a) it takes to long to solve for and i get lost

OpenStudy (amistre64):

focus!! :)

OpenStudy (anonymous):

trying nd thanks :)

OpenStudy (anonymous):

the answer u gave.. what number does that solve for?

OpenStudy (amistre64):

first one i think, dpalnc showed how to get the t coordinate, and the h coordinate is just the solution to that

OpenStudy (anonymous):

okay.

OpenStudy (amistre64):

and i think b as well by default

OpenStudy (anonymous):

can u guys come back and help